### v1

# A Look Into News Avoidance Through AWRS: An Avoidance-Aware Recommender System

Abstract.
In recent years, journalists have expressed concerns about the increasing trend of news article avoidance, especially within specific domains. This issue has been exacerbated by the rise of recommender systems. Our research indicates that recommender systems should consider avoidance as a fundamental factor. We argue that news articles can be characterized by three principal elements: exposure, relevance, and avoidance, all of which are closely interconnected. To address these challenges, we introduce AWRS, an Avoidance-Aware Recommender System. This framework incorporates avoidance awareness when recommending news, based on the premise that news article avoidance conveys significant information about user preferences. Evaluation results on three news datasets in different languages (English, Norwegian, and Japanese) demonstrate that our method outperforms existing approaches.
Recommender Systems, News Modeling, News Avoidance
1. Introduction
Recommender systems are extensively used to present users with options that align closely with their interests. In the news domain, these systems face unique challenges not typically encountered in other areas, such as the importance of timeliness, novelty, and relevance. These factors create a rapidly changing environment where news consumption patterns shift quickly.
Traditional methods relying solely on user click history and basic profiling have given way to sophisticated techniques that delve deeper into understanding user preferences and behavioral patterns. For instance, the NRMS model (Wu et al., 2019d) uses multi-head self-attention to learn news representations from news titles by modeling the interactions between words. In the user encoder, they learn representations of users from their browsed news and use multi-head self-attention to capture the relatedness between the news. Moreover, the NAML (Wu et al., 2019b) model proposes a neural news recommendation approach which can learn informative representations of users and news by exploiting different kinds of news information such as titles, bodies, and topic categories. Additionally, the LSTUR model (An et al., 2019) learns representations of news from their titles and topic categories and uses an attention network to select important words. In the user encoder, they propose to learn long-term user representations from the embeddings of their IDs.
More recently, other models have tried to increase their performance by focusing on different paradigms. For instance, the GLORY model (Yang et al., 2023) integrates global information with local user interactions to optimize content personalization. Moreover, LANCER (Bae et al., 2023) incorporates the concept of news lifetime to strategically enhance the negative sample space by perceiving that news articles have a finite influence period. Addressing challenges such as the cold-start problem and popularity bias, the PP-Rec model (Qi et al., 2021) utilizes news popularity metrics to enhance recommendation accuracy. In their method, the ranking score for recommending a candidate news to a target user is the combination of a personalized matching score and a news popularity score. Meanwhile, the CAUM model (Qi et al., 2022) leverages candidate-aware self-attention networks to capture global user interests based on candidate news items. They propose a candidate-aware CNN network to incorporate candidate news into local behavior context modeling and learn candidate-aware short-term user interest. Moreover, MANNeR (Iana et al., 2024), a modular framework for flexible multi-aspect (neural) news recommendation that supports ad-hoc customization over individual aspects at inference time, focuses on balancing recommendation performance with diversity across various metrics.
However, one factor that has become increasingly common in today’s digital landscape is the phenomenon of news avoidance. This behavior reflects a deliberate rejection or unintentional neglect of traditional news consumption, often influenced by a preference for alternative media sources (Schrøder, 2019; Villi et al., 2022; Fitzpatrick, 2022). This trend, which can be temporary or selective, underscores a nuanced understanding of how consumers interact with news, shaped by their interests and a growing skepticism towards certain topics. As the demand rises for news that aligns with personal values, especially in a competitive and diverse media landscape, the concept of relevance becomes paramount. Challenges exacerbated by the pandemic, such as news fatigue and distrust in mainstream media, emphasize the importance of recommender systems capable of recognizing and integrating avoidance behaviors.
Despite the aforementioned advancements, the integration of news avoidance behaviors into recommender systems has been overlooked, at least to the best of our knowledge. By engaging domain experts such as journalists and addressing their concerns (Schrøder, 2019; Villi et al., 2022; Fitzpatrick, 2022; Heitz et al., 2022), we explored how to perceive and integrate avoidance strategies within these news recommendation systems. This approach not only enhances the relevance and effectiveness of news recommendations but also acknowledges the evolving dynamics of user engagement with news content in today’s digital landscape.
In summary, our contributions are as follows: (1) Our model presents a novel concept, and to the best of our knowledge, we are the first to explore the perspective of avoidance in recommender systems. (2) We introduce the AWRS framework (Avoidance-Aware Recommender System), which incorporates avoidance awareness, including time and relevance modules, to improve performance in user matching recommendations. (3) Extensive experiments on three diverse real-world datasets show that AWRS consistently delivers superior performance across a wide range of metrics.
2. Related Work
2.1. Personalized News Recommender Systems
Neural content-based models have become the leading method for personalized news recommendation, surpassing traditional systems that relied on manual feature engineering (Wu et al., 2023). These Neural News Recommender (NNR) models typically include a news encoder that transforms input features into embeddings and a user encoder that creates user-level representations from clicked news (Wu et al., 2023, 2019d, 2019c; Okura et al., 2017; An et al., 2019; Wu et al., 2022b). The recommendation score is computed by comparing the candidate news embedding against the user embedding (Wang et al., 2018; Wu et al., 2019b). These models are trained using point-wise classification objectives with negative sampling (Huang et al., 2013; Wu et al., 2019d).
Personalized NNR systems focus on matching news content to user preferences, often limiting exposure to diverse viewpoints and creating ”filter bubbles” (Iana et al., 2024; Heitz et al., 2022; Pariser, 2011; Li and Wang, 2019). This leads to homogeneous news consumption, reinforcing users’ initial stances. To address these issues, researchers have explored various methods to diversify recommendations. One approach involves re-ranking personalized recommendation sets (Gharahighehi and Vens, 2021) to increase diversity. Another approach involves multi-task training of NNRs (Moreira et al., 2019; Qi et al., 2022; Wu et al., 2022a; Choi et al., 2022), where the primary personalization objective is coupled with auxiliary objectives that promote aspect-based diversification.
Recent advancements in news recommenders aim to enhance content-based personalization by considering additional aspects: (1) Sentiment - SentiRec (Wu et al., 2020b) recommends news with diverse sentiments. (2) Popularity - PP-Rec (Qi et al., 2021) incorporates news popularity to address cold-start and diversity issues. (3) Global Representations - GLORY (Yang et al., 2023) uses a global news graph and gated graph neural networks to enrich news representations. (4) Recency - LANCER (Bae et al., 2023) introduces the concept of lifetime to improve news recommendation models.
2.2. Open Possibilities and Current Limitations
As previously mentioned, news avoidance has become a growing concern for news platforms, particularly those utilizing recommender systems due to the creation of ”filter bubbles” (Pariser, 2011). We decided to investigate how avoidance affects the behavior of news article clicks. Subsequent sections will illustrate the intrinsic relationship between avoidance, exposure, and the number of clicks a news article receives. This aspect, overlooked in prior studies to the best of our knowledge, can enhance news recommendations by providing insights into the interaction between exposure and avoidance, especially regarding popular items.
In the news domain, the item with the most clicks is often highly exposed and less avoided. We argue that news article avoidance conveys significant information about user preferences. For instance, during an election, election-related articles may be popular but often avoided by those fatigued by political news. If a user reads a political article that is largely avoided by others, it suggests a strong interest in that topic. Conversely, clicking on a popular football article, one that is little avoided, might indicate less about the user’s preference and more about its current popularity. In other words, largely avoided news articles provide more information about user tastes than less avoided ones. This understanding allows us to incorporate such contextual information into recommender systems, improving the understanding of user preferences. The following sections will formally define each of the concepts discussed above.
3. Principal Elements
We identified three principal elements for characterizing a news article: exposure, avoidance, and relevance. To explain these concepts please refer to the Figure 1. In the figure, the black dashed line represents the behaviors dataset, listing news articles (impressions) at each time point $t$. For simplicity, assume each positive news candidate (clicked article) has a fixed number of negative candidates. The off-white square labeled ”exposed news articles” shows all news articles viewed by users at time $t_{1}$. Imagine a news platform where at any time $t$, a certain number of news articles $\mathcal{N}_{t}=\{n_{1},n_{2},...,n_{k}\}$ and users $\mathcal{U}_{t}=\{u_{1},u_{2},...,u_{j}\}$ interact. When user $u_{1}$ views article $n_{1}$, they are ”exposed” to it and can choose to click or not. The box contains all $k$ articles viewed by all $j$ users at time $t$.
Figure 1. A schematic for the explanation of avoidance and exposure.
The light pink rectangle, called the ”impression list,” represents all news articles viewed by user $u$, both positive and negative. An impression log records the articles displayed to a user at a specific time and their click behaviors (Wu et al., 2020a). The set of impressions at time $t$ is termed exposure. Exposure is contingent upon the publication date (time) of news articles before they can be displayed to users. Therefore, at a specific time $t$, certain news articles may be exposed while others are not, simply because they have not yet been published. This concept is illustrated by the arrow at the bottom of the figure, depicting how the impression lists change over time. Finally, we use dark blue and dark pink to indicate whether articles have been clicked or not, respectively.
3.1. Exposure
Definition 3.1 (Number of Exposures - $n_{E}(n,t)$).
The total number of times a specific news article $n$ is presented to users on the platform up to a defined time $t$. This metric quantifies all instances where the article $n$ was displayed to active users, thus providing opportunities for being clicked.
Definition 3.2 (Number of Impressions - $n_{I}(t)$).
An Impression is defined as an individual data record in a news dataset in a given time $t$.
Definition 3.3 (Exposure Per Impression - $EPI(n,t)$).
The Exposure Per Impression ($EPI$) for a news article $n$ at time $t$ is calculated as follows:
(1) $EPI(n,t)=\frac{n_{E}(n,t)}{n_{I}(t)}$
This ratio provides a relative measure of how extensively a news article $n$ has been exposed in relation to the total number of impressions recorded.
Example: At time $t_{1}$, we know that we have 100 impression lists, and the news article $n_{174}$ has been displayed in 50 of them. Thus, we have $n_{E}(n_{174},t_{1})=50$ and $n_{I}(t_{1})=100$. Hence, the exposure per impression of this news article $n_{174}$ is $EPI(n_{174},t_{1})=\frac{50}{100}=0.5$.
Figure 2. $Av(n,t)\text{ vs. }EPI(n,t)$ for MIND-small.
3.2. Avoidance
Avoidance reflects the deliberate non-engagement with specific news content by readers. This phenomenon may arise from various factors, such as a lack of interest in the topic, discomfort with the content, or distrust towards the source. Understanding the patterns of avoidance is crucial for analyzing user behavior and optimizing content delivery to enhance reader engagement.
Definition 3.4 (Avoidance - $Av(n,t)$).
Avoidance for a news article $n$ over a timeframe $t$ quantifies the proportion of potential exposures that did not result in interaction. This is mathematically defined as follows:
(2) $Av(n,t)=1-\frac{n_{clk}(n,t)}{n_{E}(n,t)}$
here, $n_{clk}(n,t)$ represents the number of clicks received by the article $n$ up to the specified time $t$, and $n_{E}(n,t)$ denotes the total number of exposures of the article $n$ within the same timeframe. The value of $Av(n,t)$ ranges from 0 to 1, where 1 indicates total avoidance (no clicks relative to exposures) and 0 signifies complete engagement (every exposure resulted in a click).
Example: Given that the news article $n_{174}$ at time $t_{1}$ received 20 clicks out of 50 exposures, its avoidance is calculated as $Av(n_{174},t_{1})=1-\frac{n_{clk}(n_{174},t_{1})}{n_{E}(n_{174},t_{1})}=1-\frac{20}{50}=0.6$.
3.2.1. Visualization of Avoidance
We analyzed the correlation among article avoidance, exposure, and the number of clicks for news articles. We plotted $Av(n,t)$ versus $EPI(n,t)$ for the MIND-small dataset on November 9, 2019, at 16:00 and 22:00, as shown in Figure 2. Additionally, we incorporated a color scale to indicate the normalized number of clicks and represented the radius of each news article based on its number of clicks. The graph demonstrates that articles maintaining a balanced level of exposure and low avoidance tend to attract more clicks over time. Articles that show an increase in $EPI$ and a decrease in $Av$ (as indicated by the red arrow) receive the highest number of clicks. For example, article N41881 maintains a stable balance between $EPI$ and $Av$ from 16:00 to 22:00, making it the most clicked news article. In contrast, article N1034 shows a shift in its $EPI$ and $Av$ balance during the same time window. It demonstrates an upward trend in exposure ($\uparrow EPI$) and an increase in avoidance ($\rightarrow Av$) over time, impacting its click rate, which indicates a gradual increase in clicks (larger circle radius). This analysis underscores how avoidance and exposure dynamics influence click rates, offering insights to enhance recommender systems.
(a) MIND-small - English
(b) Adressa one-week - Norwegian
(c) Nikkei - Japanese
Figure 3. The graph for $Av(n,t)$ vs $EPI(n,t)$ is computed for the MIND-small (a), Adressa one-week (b), and Nikkei (c) datasets using $D=5$, resulting in 25 distinct regions.
3.3. Relevance
Relevance indicates how well a news article aligns with readers’ interests, but measuring it is complex and requires specific data. Key factors in measuring relevance include: (1) time since publication, (2) news text embedding, (3) exposure per impression $EPI(n,t)$, (4) avoidance $Av(n,t)$, and (5) number of clicks $clk(n,t)$. Each factor is weighted within the model to evaluate relevance, offering a robust framework for assessing the alignment of a news article with user interests. More details on the calculation of relevance can be found in Section 5.3.
4. User Engagement Embedding
To measure the impact of avoidance and exposure on user interaction with a news article, we calculate a metric called user engagement ($\bm{ue}$)—how users interact with a specific news article based on its avoidance and exposure per impression. We divide our graph of $Av(n,t)\text{ vs. }EPI(n,t)$ into $D\times D$ distinct regions. Depending on the values of $Av(n,t)$ and $EPI(n,t)$, a news article falls into different regions. Formally, the user engagement indices are computed as:
(3) $i_{\text{ue}}=D\cdot epi_{\text{idx}}+av_{\text{idx}}$
where $epi_{\text{idx}}$ is the vector of exposure per impression indices for historical clicks, $av_{\text{idx}}$ is the vector of avoidance indices for historical clicks, and $D$ is a constant representing the number of distinct values. The user engagement embedding is then obtained using an embedding layer:
(4) $\bm{ue}=\mathbf{W}_{\text{ue}}(i_{\text{ue}})$
where $\mathbf{W}_{\text{ue}}$ is an embedding layer that maps each index to an embedding vector, and $i_{\text{ue}}$ are the computed indices from the above equation 3. The user engagement embedding ($\bm{ue}$) captures the spatial representation based on the quantized values of avoidance and exposure per impression. This embedding is used as input for our models to capture the information conveyed when a specific news article is largely avoided by many users but still clicked by a particular user. The concept behind the user engagement embedding ($\mathbf{ue}$) is to integrate avoidance-awareness into recommender models. Figure 3 illustrates this division for $D=5$, resulting in $D\times D=25$ distinct regions.
5. Methodology
As highlighted in section 2.1, different models focus on various aspects of news recommendations. We introduce a new aspect: Avoidance. Our method, AWRS, improves recommendations by incorporating avoidance-awareness. Unlike traditional methods based only on clicks and popularity, AWRS integrates avoidance and relevance. The conceptual framework is shown in Figure 4.
5.1. Problem Formulation
Given a user $u$ and a candidate news article $n_{c}$, our goal is to compute an interest score $Int_{s}$ that quantifies user $u$’s potential engagement with $n_{c}$. We evaluate a collection of candidate news articles $\mathbf{N}_{c}=[n_{c1},n_{c2},\ldots,n_{cL}]$ and recommend the highest-ranking articles to user $u$. User $u$ has a history of clicked news articles $\mathbf{H}_{u}=[h_{1},h_{2},\ldots,h_{M}]$. Each news article is characterized by its title ($\mathcal{T}$), abstract ($\mathcal{A}$), category ($\mathcal{C}_{cat}$), and associated entities ($\mathbf{E}_{i}=[e_{1},e_{2},\ldots,e_{k}]$).
Figure 4. Overview of our proposed approach for AWRS.
5.2. Document Encoder
Our news encoder, inspired by (Qi et al., 2021, 2022; Wu et al., 2019b), extracts detailed information from text and entities within a news article using a multi-head attention architecture. It incorporates embeddings from the title, category, and entities, generating embeddings with Pretrained Language Models (PLMs) following (Iana et al., 2024). We used RoBERTa Base (Liu et al., 2019) for the MIND-small dataset, NB-BERT Base (Kummervold et al., 2021) for the Adressa one-week dataset, and a finetuned version of the DeBERTa model (He et al., 2021) called Japanese DeBERTa V2111https://huggingface.co/ku-nlp/deberta-v2-base-japanese. For all models, only the last four PLM layers were fine-tuned. The final news embedding n combines $\textbf{n}_{t}$ (title), $\textbf{n}_{c}$ (category), and $\textbf{n}_{e}$ (entity). For the Nikkei and Adressa one-week datasets, only title and category information were used.
5.3. Avoidance-aware Relevance Predictor
Figure 5. (a) Avoidance-aware Relevance Predictor and (b) Avoidance-aware User Encoder Schematics.
As discussed in Section 3.3, measuring relevance at a given time $t$ involves considering several factors: the content of the article, the time elapsed since publication, avoidance and exposure per impression values, and the number of clicks. Therefore, the Avoidance-aware Relevance Predictor, illustrated in Figure 5 (a), evaluates the relevance of each previously clicked news article $n$. This assessment incorporates the time elapsed, encoded using the Time2Vec (Kazemi et al., 2019) model, resulting in the time elapsed embedding $\bm{t}_{el}$. The news embedding $\bm{n}$, obtained from the news encoder described in Section 5.2, along with the number of clicks at time $t$ denoted as $n_{clk}(n,t)$, and the impact of avoidance and exposure per impression are integrated into the user engagement embedding $\bm{ue}$, as detailed in Section 4. Consequently, the avoidance-aware relevance scores ($\textbf{r}_{aw}$) for each news article are computed as follows:
(5) $\textbf{r}_{aw}=\sigma\left(n_{\textit{clk}}(n,t)\cdot w_{\text{ctr}}+\hat{r}\cdot w_{\hat{r}}\right)$
here, $\hat{r}$ is the weighted sum of the relevance scores influenced by the news article conveyed information, defined as
(6) $\hat{r}=W\cdot\hat{r}_{\text{ic}}+(1-W)\cdot\hat{r}_{\text{tue}}$
here the weight $W$ is computed as $W=\sigma(\Psi_{1}([\textbf{n},\bm{ue},\bm{t}_{el}]))$, where [,,][$\cdot$,$\cdot$,$\cdot$][ ⋅ , ⋅ , ⋅ ] denotes vector concatenation, $\sigma$ is the sigmoid function, and $\Psi_{1}$ is a dense layer. The term $\hat{r}_{\text{ic}}$ represents the relevance score influenced by the news article embedding (n):
(7) $\hat{r}_{\text{ic}}=\Psi_{2}(\textbf{n}),\quad\hat{r}_{\text{tue}}=\Psi_{3}([\bm{ue},\textbf{t}_{el}])$
here the term $\hat{r}_{\text{tue}}$ represents the relevance score influenced by the combination of the user engagement embedding ($\bm{ue}$) and the time elapsed embedding ($\bm{t}_{el}$). Finally, the overall relevance score ($\bm{r}_{aw}$) is obtained by considering both the number of clicks up to time $t$ ($n_{clk}(n,t)$) and the computed relevance score ($\hat{r}$), adjusted by their respective weights ($w_{n_{clk}}$ and $w_{\hat{r}}$) and passed through a sigmoid activation function ($\sigma$) for normalization, as indicated by Equation 5.
5.4. Avoidance-aware User Encoder
The core idea of AWRS is to integrate avoidance-awareness. Leveraging the CAUM model’s architecture (Qi et al., 2022), which facilitates the extraction of how closely candidate news articles relate to previously clicked ones, AWRS enhances the candidate-aware attention network with user engagement embeddings. This enhancement refines the comprehension of user preferences by examining the categories of news articles users interact with, as well as their levels of exposure and avoidance. This approach yields a nuanced depiction of user interests.
5.4.1. Historical and Candidate News Vector Concatenation
For each historical and candidate news item, we enhance its representation by appending the corresponding user engagement embeddings. This generates enriched feature vectors for a more tailored recommendation approach. Given a set $\mathbf{N}_{c}$ of candidate news articles and a set $\mathbf{H}_{u}$ of historical clicks for a user $u$, we concatenate the user engagement embedding $\bm{ue}$ for each item. Let $\mathscr{H}$ and $\mathscr{N}$ be the sets with concatenated user engagement embeddings. Formally, for each $i$-th item $\bm{h}_{i}=[\textbf{h}_{u_{i}},\bm{ue}_{h_{i}}]$ and $\bm{n}_{i}=[\mathbf{n}_{c_{i}},\bm{ue}_{c_{i}}]$. Where $\bm{ue}_{h_{i}}$ and $\bm{ue}_{c_{i}}$ are the user engagement embeddings for the $i$-th historical and candidate news items, respectively, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes vector concatenation. Thus, we have the sets:
(8) $\mathscr{H}=[\bm{h}_{1},...,\bm{h}_{M}],\quad\mathscr{N}=[\bm{n}_{c1},...,\bm{n}_{cL}]$
5.4.2. Avoidance-aware Self-Attention Layer
Given the history of clicked news articles $\mathscr{H}_{u}^{M}$, where $M$ represents the total number of articles previously clicked by the user $u$, we use multiple self-attention heads to assess similarities and extract relatedness information between the $i$-th and $j$-th historical clicks, $\textbf{h}_{i}$ and $\textbf{h}_{j}$, respectively:
(9) $\hat{r}^{k}_{i,j}=\textbf{q}^{T}_{i}\textbf{W}^{r}_{k}\bm{h}_{j},\quad\textbf{q}^{T}_{i}=\textbf{Q}_{u}\bm{h}_{i}$
here, $\hat{r}^{k}_{i,j}$ is the attention score from the $k$-th head, with $\textbf{Q}_{u}$ as the projection matrix and $\textbf{W}^{r}_{k}$ as the parameters of the $k$-th attention head. We adaptively select significant long-range relatedness metrics to model user interest in the candidate news $\mathscr{N}$ based on their contextual relevance
(10) $r^{k}_{i,j}=\hat{r}^{k}_{i,j}+\textbf{q}^{T}_{c}\textbf{W}^{r}_{k}\bm{h}_{j}$
where $\textbf{q}^{T}_{c}=\textbf{Q}_{c}\mathscr{N}$. This enhances the attention score with candidate-specific adjustments. The augmented representation $\textbf{l}^{k}_{i}$ for each click is formulated through attention weights:
(11) $\textbf{l}^{k}_{i}=\textbf{W}^{k}_{o}\sum_{j=1}^{N}\gamma^{k}_{j}\bm{h}_{j},\quad\gamma^{k}_{j}=\frac{\exp(r^{k}_{i,j})}{\sum_{p=1}^{N}\exp(r^{k}_{i,p})}$
$\textbf{W}^{k}_{o}$ is the projection matrix for the $k$-th head. The comprehensive contextual representation $\textbf{l}_{i}$ for each click is derived by merging outputs from all $K$ attention heads. Note that by incorporating user engagement embeddings, our relatedness information now considers the influence of both avoidance and exposure, providing more context to user click patterns.
5.4.3. Avoidance-aware CNN Layer
The historical news vectors undergo processing through a convolutional neural network (CNN) equipped with self-attention mechanisms. According to (Qi et al., 2022), this involves applying multiple filters to capture patterns among local contexts of adjacent clicks and candidate news, formulated as
(12) $\textbf{s}_{i}=\textbf{W}_{awc}[\bm{h}_{i-h};\dots;\bm{h}_{i};\dots;\bm{h}_{i+h};\mathscr{N}]$
[…]
5.4.4. Avoidance-aware Final Attention Layer
We use a candidate-aware attention network to model the importance of clicked news based on their relevance to avoidance-aware candidate news $\mathscr{N}$. This builds the avoidance-aware user embedding representation $\textbf{u}_{aw}$, given by $\textbf{u}_{aw}=\sum^{N}_{i=1}\alpha_{i}\textbf{m}_{i}$ where $\alpha_{i}$ is the weight of the $i$-th click:
(13) $\alpha_{i}=\frac{\exp(\Phi_{2}(\textbf{m}_{i},\mathscr{N}))}{\sum^{N}_{j=1}\exp(\Phi_{2}(\textbf{m}_{j},\mathscr{N}))}$
here, $\Phi_{2}$ is a dense layer. This approach encodes avoidance-aware user interests relevant to the candidate news into $\textbf{u}_{aw}$, enhancing interest matching accuracy. The general overview of the Avoidance-aware User Encoder is shown in Figure 5 (b).
5.4.5. Final Relevance Scores
The relevance scores combine the news candidate’s avoidance-awareness $\mathscr{N}$ with the score from the avoidance-aware user vector $\textbf{u}_{aw}$. Specifically, the preliminary interest scores are calculated as $Int_{s}^{\prime}=\mathscr{N}^{T}\cdot\textbf{u}_{aw}$. However, to obtain the final scores, we apply our Avoidance-aware Relevance Predictor, which considers the time-varying nature of avoidance, as shown in Figure 4. Thus,
(14) $Int_{s}=(1-\eta)\cdot\mathbf{r}_{aw}+\eta\cdot Int_{s}^{\prime}$
where $\eta=\sigma(\Phi_{3}(\mathbf{u}_{aw}))$. Here, $\Phi_{3}$ is a dense layer, $\mathbf{r}_{aw}$ is the relevance avoidance-aware score as defined earlier by equation 5, and $\eta$ is computed from the user representation $\mathbf{u}_{aw}$ using a dense network with sigmoid activation. This approach allows our model to better capture patterns related to the information conveyed by theavoidance and exposure values of a news article.
5.5. Loss Function
Inspired by (Wu et al., 2019d), we use negative sampling during model training. For each clicked news item (positive sample), we randomly sample $K$ non-clicked items (negative samples) from the same impression and shuffle their order. The click probability score of the positive item is $\hat{y}^{+}$, and the scores for the $K$ negative items are $[\hat{y}^{-}_{1},\hat{y}^{-}_{2},\ldots,\hat{y}^{-}_{K}]$. Then, these scores are normalized using the softmax function given by
(15) $p_{i}=\frac{\exp(\hat{y}^{+}_{i})}{\exp(\hat{y}^{+}_{i})+\sum_{j=1}^{K}\exp(\hat{y}^{-}_{i,j})}$
We frame the click prediction task as a pseudo $(K+1)$-way classification problem. The loss function is the negative log-likelihood of all positive samples $S$: $\mathcal{L}=-\sum_{i\in S}\log(p_{i})$. This method trains the model to distinguish between clicked and non-clicked news items.
6. Experiment
6.1. Experimental Settings
6.1.1. Datasets
We evaluated AWRS on three datasets: MIND-small (English) (Wu et al., 2020a), Adressa one-week (Norwegian) (Gulla et al., 2017), and Nikkei (Japanese). Nikkei Inc. [^2] https://www.nikkei.com/, based in Japan, is a leading media corporation renowned for its comprehensive coverage of business, economic, and financial news. Table 1 summarizes the train, validation, and test splits used in our study. For Nikkei, we analyzed 15,803 news articles, dividing user click behavior into training (Jan 16-20, 2023), validation (Jan 21, 2023), and testing (Jan 22, 2023) sets. The Adressa one-week dataset includes 22,136 news articles, with user click behaviors split into training (Jan 1-5, 2017), validation (Jan 6, 2017), and testing (Jan 7, 2017) sets. And the MIND-small dataset comprises 93,698 news articles, with user click behaviors segmented into training (Nov 9-13, 2019), validation (Nov 14, 2019), and testing (Nov 15, 2019) sets.
Table 1. Dataset Statistics
Dataset Train Validation Test
Impr./Users Impr./Users Impr./Users
Adressa 181,279 / 83,599 36,412 / 27,943 145,626 / 68,565
MIND-small 124,229 / 45,214 29,498 / 19,703 70,938 / 48,593
Nikkei 137,142 / 23,139 10,560 / 6,201 9,695 / 5,805
6.1.2. Baselines
We conducted a comparative analysis of the AWRS model against several state-of-the-art (SOTA) baseline models to evaluate its performance. The models compared include: (1) NRMS (Wu et al., 2019d), (2) NAML (Wu et al., 2019b), (3) LSTUR (An et al., 2019), (4) TANR (Wu et al., 2019a), (5) SentiRec (Wu et al., 2020b), (6) MINER (Li et al., 2022), (7) MINS (Wang and Lu, 2022), (8) CenNewsRec (Qi et al., 2020), (9) MANNeR-CR [^3] For the Nikkei dataset and MANNeR-CR baseline model, we had to reduce the maximum number of tokens due to memory constraints. (Iana et al., 2024), (10) PP-REC (Qi et al., 2021), and (11) CAUM (Qi et al., 2022).
6.1.3. Evaluation Metrics
In accordance with prior research (Qi et al., 2022; Yang et al., 2023), we evaluate model performance using Area under the ROC Curve (AUC), Mean Reciprocal Rank (MRR), and normalized discounted cumulative gain (nDCG@5 and nDCG@10).
Table 2. Performance comparisson for the Nikkei, MIND-small, and Adressa one-week datasets. The best results are bolded, and the second best are underlined. All the results reported here used $D=5$.
Nikkei MIND-small Adressa one-week
MODEL AUC MRR nDCG@5 nDCG@10 AUC MRR nDCG@5 nDCG@10 AUC MRR nDCG@5 nDCG@10
NRMS 0.5140 0.3110 0.2117 0.3069 0.5000 0.2731 0.2532 0.3198 0.5847 0.2102 0.1824 0.2894
NAML 0.5011 0.3323 0.2324 0.3322 0.5000 0.2806 0.2606 0.3246 0.5000 0.2267 0.1984 0.2945
LSTUR 0.4953 0.2878 0.1929 0.2883 0.5000 0.3000 0.2830 0.3468 0.5358 0.3091 0.2931 0.3576
TANR 0.5007 0.3005 0.2059 0.3011 0.5224 0.2964 0.2773 0.3425 0.5359 0.2622 0.2491 0.3136
SentiRec 0.5000 0.3051 0.2042 0.2991 0.5289 0.2773 0.2642 0.3281 0.5313 0.2087 0.1754 0.2831
MINER 0.5439 0.3296 0.2342 0.3613 0.6141 0.2722 0.2591 0.2385 0.6012 0.2165 0.1908 0.3054
MINS 0.5007 0.2723 0.1674 0.2677 0.4266 0.1995 0.1756 0.2385 0.7018 0.3511 0.3837 0.4703
CenNewsRec 0.5249 0.2873 0.1679 0.2521 0.4837 0.1992 0.1770 0.2384 0.5980 0.3211 0.3194 0.3781
MANNeR-CR 0.5000 0.3422 0.2422 0.3451 0.5917 0.3018 0.2826 0.3469 0.5000 0.2248 0.1741 0.2481
PP-REC 0.4970 0.2731 0.1869 0.2808 0.5686 0.2912 0.2695 0.3301 0.6729 0.4335 0.4129 0.4714
CAUM 0.5606 0.3278 0.2391 0.3411 0.5528 0.3133 0.2972 0.3615 0.6300 0.2476 0.2434 0.3014
AWRS 0.6723 0.4973 0.4028 0.4781 0.5978 0.3267 0.3105 0.3722 0.7670 0.4826 0.5134 0.5627
6.1.4. Environment Configuration
During training, we used mixed precision and the Adam optimizer (Kingma and Ba, 2017). The learning rates were 1e-5 for MIND-small and Nikkei, and 1e-6 for Adressa one-week. We used $k=4$ negative samples for all datasets. Each model for the MIND-small was trained for 10 epochs and for Adressa one-week and Nikkei 3 epochs, using a single NVIDIA A100 GPU, with implementations in the newsreclib framework [^4] Upon publication, we plan to release parts of the code to the research community for further development and replication of our results. (Iana et al., 2023). When calculating avoidance and user engagement embedding, we measured it every 1 hour for the MIND-small dataset, every 5 hours for the Adressa one-week dataset, and every 2 hours for the Nikkei dataset.
7. Results
With the results at hand we comprehensively evaluated AWRS by answering the following research questions:
  • RQ1 (Accuracy): Does the proposed AWRS method outperform existing baselines in terms of accurately predicting user-clicked news articles?
  • RQ2 (Effect of Each Component): How does each component of our proposed architecture contribute to the accuracy scores? Specifically, what is the impact of adding the Avoidance-aware User Encoder and Avoidance-aware Relevance Predictor components?
  • RQ3 (Effect of Varying $D$): To compute the user engagement embedding ($\bm{ue}$), we divided our avoidance region space ($av_{idx}$) and exposure per impression ($epi_{idx}$) into distinct regions determined by the constant $D$. How does accuracy vary when changing this constant?
  • RQ4 (Performance without PLMs): How does our model perform in the absence of pretrained language models (PLMs)?
  • RQ5 (Limitations): What are some of the limitations of our model compared to existing baselines?
7.1. RQ1 (Accuracy)
Table 2 presents a performance comparison between the baselines and our proposed AWRS model. As shown in the table, our model demonstrates improvements in nearly all key metrics, attributed to its enhanced awareness of avoidance and exposure. For instance, if a user opts to read an article about chess, a typically avoided topic, it suggests a strong interest in this subject. Conversely, an article about the Oscars, which is generally less avoided, might be clicked due to its current popularity rather than genuine interest. It is crucial to note that such patterns may change over time. For example, during a specific period involving controversial news about chess, articles on chess might become less avoided, whereas Oscars articles might be more avoided if they are far from the premiere date. Therefore, we argue that avoidance and time are closely related. This understanding allows us to delve deeper into our discussion for each dataset.
7.1.1. Nikkei
Our proposed AWRS model is the best-performing model for this dataset, with the second-best performance varying across different metrics. The second highest AUC was achieved by CAUM (Qi et al., 2022), the second best MRR and nDCG@5 by MANNeR (Iana et al., 2024), and the second highest nDCG@10 by MINER (Li et al., 2022). All these baselines have complex architectures for encoding user behaviors, demonstrating their efficacy in predicting user preferences. However, by incorporating user engagement embeddings with our Avoidance-aware User Encoder alongside our Avoidance-aware Relevance Predictor as indicated in figures 4 and 5, our model achieves a better understanding of user patterns. This leads to an avoidance-aware contextualized assessment of each article, enabling our model to extract more information by understanding the avoidance of news articles.
7.1.2. MIND-small
For this dataset, our model outperforms others in terms of MRR, nDCG@5, and nDCG@10. However, it performs slightly worse than MINER (Li et al., 2022) for AUC. Generally, the second-best model was CAUM (Qi et al., 2022), for reasons similar to those elaborated for Nikkei. We argue that by fully capturing the avoidance behavior of news articles, our model better understands user patterns.
7.1.3. Adressa one-week
We employed various baselines, each with different advantages and disadvantages. The Adressa dataset contains click patterns for news articles published long ago. Despite encompassing the initial weeks of 2017, the dataset includes user interactions with very old news articles (e.g., a user clicking on a news article from the 2000s during the early weeks of 2017). Therefore, baselines such as PP-REC (Qi et al., 2021), which consider time as a key metric, perform very well on this dataset. However, our AWRS model not only incorporates time as a factor through the Time2Vec (Kazemi et al., 2019) module, but also considers avoidance, enhancing the contextualization for the model’s decision-making process.
7.2. RQ2, RQ3 and RQ4 (Ablation Studies)
We aim to address research questions RQ2, RQ3, and RQ4 by exposing an ablation study, as demonstrated in the following sections.
Table 3. Impact of the Avoidance-aware User Encoder and the Avoidance-aware Relevance Predictor modules on AWRS.
Dataset AWRS Model AUC MRR nDCG@5 nDCG@10
MIND-small only rel 0.5818 0.3244 0.3083 0.3690
only avoid 0.5838 0.3077 0.2884 0.3511
avoid + rel 0.5978 0.3267 0.3105 0.3722
Adressa one-week only rel 0.6354 0.3621 0.3948 0.4526
only avoid 0.5423 0.2564 0.2204 0.2499
avoid + rel 0.7670 0.4826 0.5134 0.5627
Nikkei only rel 0.6362 0.4638 0.3701 0.4534
only avoid 0.6522 0.4889 0.3784 0.4568
avoid + rel 0.6723 0.4973 0.3784 0.4781
Table 4. AWRS scores for different datasets with varying $D$ values.
Dataset Metrics D = 5 D = 7 D = 10 D = 15 D = 20
MIND-small AUC 0.5978 0.6213 0.6430 0.6479 0.6494
MRR 0.3267 0.3214 0.3499 0.3481 0.3409
nDCG@5 0.3105 0.3049 0.3285 0.3289 0.3226
nDCG@10 0.3722 0.3684 0.3914 0.3923 0.3864
Adressa one-week AUC 0.7670 0.6561 0.6997 0.6443 0.7652
MRR 0.4826 0.4310 0.4598 0.4109 0.5055
nDCG@5 0.5134 0.4172 0.4799 0.3869 0.5300
nDCG@10 0.5627 0.4670 0.5062 0.4459 0.6033
Nikkei AUC 0.6723 0.6466 0.6252 0.6283 0.6424
MRR 0.4973 0.4804 0.4168 0.4497 0.4494
nDCG@5 0.4028 0.3732 0.3178 0.3405 0.3435
nDCG@10 0.4781 0.4531 0.4096 0.4279 0.4330
Table 5. Performance comparison for the MIND-small dataset. All the results reported here used $D=5$ and GloVe.
Model AUC MRR nDCG@5 nDCG@10
SentiRec 0.5425 0.3060 0.2826 0.3486
TANR 0.5389 0.3240 0.3039 0.3683
LSTURINI 0.5482 0.3194 0.2995 0.3644
NAML 0.5010 0.3350 0.3176 0.3802
NRMS 0.5571 0.2924 0.2730 0.3370
GLORY 0.6603 0.2978 0.3218 0.3825
CAUM 0.6200 0.3414 0.3207 0.3854
AWRS 0.6613 0.3425 0.3249 0.3896
7.2.1. RQ2 (Effect of Each Component)
To assess the influence of each component in our architecture, specifically the Avoidance-aware User Encoder and the Avoidance-aware Relevance Predictor, we conducted experiments by removing each component separately and calculating the MRR, nDCG@5, and nDCG@10 scores for the MIND-small, Adressa one-week, and Nikkei datasets. As shown in Table 3, it is evident that each component individually underperforms compared to when both components are used together. We argue that this is because the Avoidance-aware User Encoder is responsible for understanding how user behavior historically patterns in terms of avoidance. It analyzes how the news articles that the user has clicked on were avoided. For example, is the user more interested in less avoided or more avoided news articles? Conversely, the Avoidance-aware Relevance Predictor gains a better understanding of the temporal and global influence of avoidance. Thus the impact of avoidance on the decision of which news article is more relevant can vary depending on the timing of the recommendation. This dual approach enhances the model’s ability to accurately predict user preferences.
7.2.2. RQ3 (Effect of Varying $D$)
To investigate the impact of adjusting the parameter $D$ on the AWRS model, we conducted experiments with values $D=5,7,10,15,$ and $20$. As detailed in Table 4, varying $D$ has a notable effect on model performance, albeit with nuanced outcomes. Notably, optimal scores for the Nikkei dataset are observed at $D=5$ and $D=7$, whereas for the Adressa one-week dataset, peak performance is achieved at $D=20$. In the case of the MIND-small dataset, the highest AUC score emerges at $D=20$, with MRR peaking at $D=10$, and both nDCG@5 and nDCG@10 showing optimal results at $D=15$. Our findings suggest that adjusting $D$ influences the embedding space’s complexity, potentially leading to overfitting or underfitting of avoidance behavior in news article recommendations. Thus, our grid search aimed to pinpoint the most effective $D$ values tailored to each dataset and metric.
7.2.3. RQ4 (Performance without PLMs)
We investigated how excluding pretrained language models (PLMs) in favor of GloVe embeddings (Pennington et al., 2014) would affect our model’s performance. For evaluation, we utilized the same datasplit from the MIND-small dataset described in Section 6.1.1. In this comparison, we included several baseline models for a more comprehensive evaluation: we employed GLORY (Yang et al., 2023), SentiRec (Wu et al., 2020b), CAUM (Qi et al., 2022), TANR (Wu et al., 2019a), NRMS (Wu et al., 2019d), NAML (Wu et al., 2019b), and LSTUR (An et al., 2019). Table 5 demonstrates that our model, AWRS, exhibits superior performance compared to such models. The reduced performance gains compared to PLMs are due to PLMs’ advanced capability in capturing more detailed contextual information from news articles.
7.3. RQ5 (Limitations)
From our exploration of current methodologies in news article recommendation systems, we identify two notable approaches. Firstly, there are methods like PP-REC (Qi et al., 2021) that leverage contextual information such as time to enhance recommendations. Secondly, methods such as those discussed in GLORY (Yang et al., 2023) utilize global information derived from static global graphs. Our proposed model aligns with the former category, leveraging time-varying avoidance data of news articles to enhance user recommendations. However, like all models, our approach has its drawbacks. For instance, one limitation of our model is its increased memory requirements due to the need for extensive per-unit-time information such as avoidance ($Av$) and Exposure Per Impression ($EPI$). Despite this increased memory demand, we did not observe a corresponding increase in computational training time for our models.
8. Conclusion
In this study, we emphasize the importance of avoidance awareness. Our research highlights that integrating avoidance behaviors can significantly enhance the predictive accuracy of recommender systems. Essentially, analyzing the timely avoidance behavior of news articles provides deeper insights into user preferences. This understanding enables us to incorporate contextual information into recommender systems, thereby refining our understanding of user preferences. By integrating our Avoidance-aware Relevance Predictor and Avoidance-aware User Encoder modules, we have developed a novel model named AWRS. This model demonstrates robust performance across three real-world datasets, showing improvements in various benchmark metrics. Our approach not only shows enhanced accuracy but also establishes a foundation for future research into understanding user behavior through avoidance.
Acknowledgements.
This work is partially supported by ”Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures (JHPCN)” in Japan (Project ID: jh241004), JSPS KAKENHI Grant Number 23K28098, and the Monbukagakusho: MEXT (Ministry of Education, Culture, Sports, Science and Technology - Japan) scholarship.